Systems Staff Meeting Notes 11/17/03

Systems Staff Meeting Notes
November 17, 2003

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Hilton Soto
Announcements
  1. Gators eeked out another win – this time over South Carolina.
  2. The power is up in the new addition of Bobbie’s house.
Moves/Construction Status
  1. Suzy has moved to DLC. Her phone extension should stay the same.
  2. Transition Monitoring Meeting – status of West – snags: the grad carrels need revamping. Dale wanted it redrawn. Groundbreaking still scheduled for December.
  3. Electricians are busy at BDF – ‘on schedule’, per Rob Roberts. We have not heard anything else from CNS about how they are going to get networking out there.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Rush to order four laser printers so the Teak Suite exiles will be able to print. They now have a total of 10 – each Director plus the Personnel staff.
  2. Web access to email is broken. Unknown if it is due to the OfficeScan updates. Unknown when/if it can be fixed. POP or IMAP client can still be used to access mail from home.
  3. Hang was trying to upgrade his OfficeScan yesterday – unsuccessfully. Seems that OfficeScan is not installing correctly. May redirect it to NT25 – the registry shows server by IP and name. Export keys wouldn’t have to go through the registry editor. May move the whole thing to NT14 – need to test.
  4. Friday – Documents machine was taken over – it was sending out Prozac ads – the only symptom the machine exhibited was sluggishness due to constant network activity. This type of problem is only getting worse.
Liaisons
  1. Shuffling of Liaisons – as a rule, when people from West move, their liaisons are still responsible for them. However, this is not the case for the Collection Management people (David Hickey, Peter Malanchuk, Dan Reboussin, and the Africana grad assistant) who have moved into Special Collections. They may be a permanent move – currently all report to Shaddy, so Mil/Flo are their liaisons.
Networking
  1. Power poles are currently being installed in Room 1 East.
  2. Work on Room 201B (janitor) is underway.
  3. Tim from CNS should be in contact soon.
  4. MSL107 is out on bid.
Software Projects Status
  1. Cynthia divided out SysRequest into 3 categories: HW/SW, Programming Requests and Systems Service Requests.
  2. Arc java servlet – ready to switch to TomCat.
  3. FRDI – Hilton is making some improvements then he’ll demo.
  4. Personnel – Web monthly OPS report – needs revamping to show the history of OPS budgets across the years.
Web Projects
  1. One more meeting between Tom, Martha, Suzy and Rich - then Tom should be ready to take over Suzy’s Web duties.


Back to the Systems Homepage

Last updated November 24, 2003
by Debra Fetzer
debharr@mail.uflib.ufl.edu